# Context-free graphs and their transition groups

### v2

@@ 9.1. Context-free graphs and free product
In this section we prove that the free product of directed graphs preserves the property of being context-free. For this purpose, we need to slightly adapt the notion of isomorphism and end-cone of an $A_{1}$-digraph to include the notion of vertex coloring. A vertex-coloring of a graph $\Gamma$ is just a map $\chi:V(\Gamma)\to C$ onto some set of colors $C$. An isomorphism $\psi:\Gamma^{(1)}\to\Gamma^{(2)}$ between two $A$-digraphs with color maps $\chi_{1}:V(\Gamma)\to C$, $\chi_{2}:V(\Gamma^{(2)})\to C$ is an isomorphism of $A$-digraphs that preserves the colorings: $\chi_{1}=\chi_{2}\circ\psi$. For technical reasons that will be clear later, we call a coloring of a rooted $A$-digraph $(\Gamma,\gamma_{0})$ any map $\chi:V(\Gamma)\setminus\{\gamma_{0}\}\to C$, and we use the notation $(\Gamma,\gamma_{0};\chi)$ to stress the coloring function. For such colored rooted $A$-digraph $(\Gamma,\gamma_{0};\chi)$, we may regard an end-cone $\Gamma(v,\gamma_{0})$ as a colored $A$-digraph with coloring map $\chi\mid_{V(\Gamma(v,\gamma_{0}))}$ given by the restriction of $\chi$ to the set of vertices of the end-cone $\Gamma(v,\gamma_{0})$. We therefore say that the colored $A$-digraph $(\Gamma,\gamma_{0};\chi)$ is context-free if there are finitely many end-cones up to end-isomorphism (in the aforementioned sense). We stick to the notation introduced in Definition 9.1, and let us suppose that $\Theta_{1}$ ($\Theta_{2}$) is a collection of $A_{1}$-digraphs ($A_{2}$-digraphs) with $A_{1}\cap A_{2}=\emptyset$ and with gluing maps $\Psi$. We say that $\Gamma^{(j)}\in\Theta_{1}$ ($\Lambda^{(j)}\in\Theta_{2}$) is context-free if the colored rooted $A_{1}$-digraph $(\Gamma^{(j)},\gamma_{0}^{j},\psi^{1}_{j})$ ($A_{2}$-digraph $(\Lambda^{(j)},\lambda_{0}^{j},\psi^{2}_{j})$) is context-free. If all the colored $A_{i}$-digraphs of $\Theta_{1}\cup\Theta_{2}$ are context-free, we say that $(\Theta_{1},\Theta_{2},\Psi)$ is a context-free ensemble. We have the following theorem.
@@
For a context-free ensemble $(\Theta_{1},\Theta_{2},\Psi)$, the free-product $\Theta_{1}\star_{\Psi}\Theta_{2}$ is a context-free $A_{1}\cup A_{2}$-digraph.
@@
an then by identifying $z$ with $\psi^{2}_{s}(z)$. We prove that we have finitely many end-cones up to end-isomorphism (in the usual sense). Now, let $\Omega(v^{*}\gamma_{i}^{j}\lambda_{\ell}^{s},\alpha)$ be another end-cone for which there is an end-isomorphism $\varphi:\Lambda^{s}(\lambda^{s}_{t},\lambda^{s}_{0})\to\Lambda^{s}(\lambda^{s}_{\ell},\lambda^{s}_{0})$ of end-cones of the same $A_{2}$-digraph $\Lambda^{(s)}$ that preserves the coloring. Since the coloring map $\psi^{2}_{s}$ is commuting with $\varphi$, for any $z\in V(\Lambda^{s}(\lambda^{s}_{t},\lambda^{s}_{0}))$ we have that $\psi^{2}_{s}(z)=\psi^{2}_{s}(\varphi(z))=\gamma_{0}^{h}$, so both $z$ and $\varphi(z)$ have the same color, hence in both the vertex $z$ and $\varphi(z)$ we are gluing the same digraph $\Gamma^{(h)}(\uparrow_{\Psi})$. Therefore, it is straightforward to show that $\varphi$ may be extended
@@
by locally extending $\varphi$ from each disjoint copy $\Gamma^{(h)}(\uparrow_{\Psi})$ associated to $z$, to the same digraph $\Gamma^{(h)}(\uparrow_{\Psi})$ associated to $\varphi(z)$. Now, since we are performing the identification $z=\psi^{2}_{s}(z)$ in the domain of $\widehat{\varphi}$ and this is the same as identifying $\varphi(z)$ with $\psi^{2}_{s}(\varphi(z))$, we have that $\widehat{\varphi}$ induces an isomorphism $\phi:\Omega(v^{*}\gamma_{i}^{j}\lambda_{t}^{s},\alpha)\to\Omega(v^{*}\gamma_{i}^{j}\lambda_{\ell}^{s},\alpha)$. Note that the frontier vertices of $\Omega(v^{*}\gamma_{i}^{j}\lambda_{t}^{s},\alpha)$ are (up to renaming) the same as $\Lambda^{s}(\lambda^{s}_{t},\lambda^{s}_{0}))$, and similarly for $\Omega(v^{*}\gamma_{i}^{j}\lambda_{\ell}^{s},\alpha)$. Hence, $\phi$ is mapping the frontier vertices of $\Omega(v^{*}\gamma_{i}^{j}\lambda_{t}^{s},\alpha)$ into the frontier vertices of $\Omega(v^{*}\gamma_{i}^{j}\lambda_{\ell}^{s},\alpha)$, hence they are end-isomorphic. We have just shown that the end-isomorphism classes of the end-cones of $\Omega$ of vertices colored by $2$ depends on the end-isomorphism (preserving the coloring) classes of the digraphs $\{\Lambda^{(1)},\ldots,\Lambda^{(m)}\}$ of $\Theta_{2}$, which are finitely many. Similarly, one may show the same for the end-cones of the vertices colored by $1$, and this concludes the proof of the theorem. ∎
Question 9.6.
It would be interesting to see what is the corresponding construction from the machine point of view. More precisely, if one considers the pushdown automata $M_{i}$ that accepts the language of the words labelling a circuit at each root of each context-free graph of $\Theta_{1}\cup\Theta_{2}$, what is the construction, involving each $M_{i}$, that generates the pushdown-automaton accepting the language of circuit of the digraph $\Theta_{1}\star_{\Psi}\Theta_{2}$?
@@
Figure 5. On the upper part the digraphs $\Theta_{1},\Theta_{2}$, in the lower picture the antenna digraph $(\Theta_{1}\star_{\Psi}\Theta_{2},\gamma)$.
10. Some examples using the free product of inverse graphs
@@
In this section we use the free product of context-free graphs to construct some examples of interesting CF-TR groups. In the first example, the antenna digraph, we show that in general transition groups of context-free graphs are not residually-finite. In the second one, the comb digraph, we prove that the associated transition group is a co-context-free group that is not poly-context-free, showing that these classes of groups are different. The third example addresses the question about the possibility of having arbitrarily large finite subgroups, we answer positively to this by providing a infinite torsion free subgroup.
Definition 10.1 (The Antenna digraph).
Let $A_{1}=\{c,c^{-1}\}$, $A_{2}=\{a,a^{-1},b,b^{-1}\}$. Consider the rooted inverse $A_{1}$-digraph $(\Gamma,\gamma)$ which is the line
@@
$\gamma_{i}\smash{\stackrel{{\scriptstyle\text{\tiny{$c$}}}}{{\relbar\joinrel\longrightarrow}}}\gamma_{i+1},\;\gamma_{i+1}\smash{\stackrel{{\scriptstyle\text{\tiny{$c^{-1}$}}}}{{\relbar\joinrel\longrightarrow}}}\gamma_{i},\quad\forall i\in\mathbb{Z},\mbox{ where the root }\gamma=\gamma_{0}$
@@
and the two rooted inverse $A_{2}$-digraphs $(\Lambda_{a},\lambda_{a}),(\Lambda_{b},\lambda_{b})$ which are lines with loops defined by
@@
$\lambda_{a,i}\smash{\stackrel{{\scriptstyle\text{\tiny{$b$}}}}{{\relbar\joinrel\longrightarrow}}}\lambda_{a,i},\;\lambda_{a,i}\smash{\stackrel{{\scriptstyle\text{\tiny{$b^{-1}$}}}}{{\relbar\joinrel\longrightarrow}}}\lambda_{a,i},\,\lambda_{a,i}\smash{\stackrel{{\scriptstyle\text{\tiny{$a$}}}}{{\relbar\joinrel\longrightarrow}}}\lambda_{a,i+1},\;\lambda_{a,i+1}\smash{\stackrel{{\scriptstyle\text{\tiny{$a^{-1}$}}}}{{\relbar\joinrel\longrightarrow}}}\lambda_{a,i},\quad\forall i\in\mathbb{Z},\,\,\lambda_{a,0}=\lambda_{a}$
@@
$\lambda_{b,i}\smash{\stackrel{{\scriptstyle\text{\tiny{$a$}}}}{{\relbar\joinrel\longrightarrow}}}\lambda_{b,i},\;\lambda_{b,i}\smash{\stackrel{{\scriptstyle\text{\tiny{$a^{-1}$}}}}{{\relbar\joinrel\longrightarrow}}}\lambda_{b,i},\,\lambda_{b,i}\smash{\stackrel{{\scriptstyle\text{\tiny{$b$}}}}{{\relbar\joinrel\longrightarrow}}}\lambda_{b,i+1},\;\lambda_{b,i+1}\smash{\stackrel{{\scriptstyle\text{\tiny{$b^{-1}$}}}}{{\relbar\joinrel\longrightarrow}}}\lambda_{b,i},\quad\forall i\in\mathbb{Z},\,\,\lambda_{b,0}=\lambda_{b}$
@@
that are constant maps. By putting $\Theta_{1}=\{(\Gamma,\gamma)\}$ and $\Theta_{2}=\{(\Lambda_{a},\lambda_{a}),(\Lambda_{b},\lambda_{b})\}$, the rooted inverse $A_{1}\cup A_{2}$-digraph $(\mathfrak{A},\gamma)=(\Theta_{1}\star_{\Psi}\Theta_{2},\gamma)$ is called the antenna digraph, see Fig. 5. Moreover, it is easy to see that $(\Theta_{1},\Theta_{2},\Psi)$ is a context-free-ensemble, thus by Theorem 9.5 $(\mathfrak{A},\gamma)$ is a context-free inverse graph. Note that with the convention of the free product that we have given, the root $\gamma$ of $\Theta_{1}\star_{\Psi}\Theta_{2}$ is identified with $\lambda_{b}$.
@@
(11) $wc^{s}\cdot b^{t}=\begin{cases}wc^{s},&\mbox{ if }s>0\\ wc^{s}b^{t},&\mbox{ otherwise }\end{cases},\,\,wc^{s}\cdot a^{t}=\begin{cases}wc^{s}a^{t},&\mbox{ if }s\geq 0\\ wc^{s},&\mbox{ otherwise }\end{cases}\,\,$
@@
        Now, if $t\neq-k$, then $v^{\prime}c^{-\ell}b^{t}\cdot(b^{k}c^{-\alpha}a^{-s}c^{\alpha}b^{-k}c^{-\alpha})=v^{\prime}c^{-\ell}b^{t+k}c^{-\alpha}\cdot(a^{-s}c^{\alpha}b^{-k}c^{-\alpha})=v^{\prime}c^{-\ell}b^{t}c^{-\alpha}=v$. Otherwise, suppose that $t=-k$. In this case we have
@@
  • the last case is when $x=a$. In this case we have to subdivide the following two sub-cases:
@@
The group $G=\mathcal{G}(\mathfrak{A})$ defined by the antenna inverse graph $\mathfrak{A}$ is not residually finite.
@@
In light of Theorem 6.2 and the preceding theorem, we observe that, if Brough Conjecture 1 regarding poly-context-free groups is true, then the class of all poly-context-free groups would be strictly contained within the class CF-TR. Although we are not able to show that the class of poly-context-free is included into our class, in the next example we show that these two classes are different. This is done by defining the following context-free digraph.
Figure 6. On the upper part the digraphs $\Theta_{1},\Theta_{2}$, in the lower picture the comb digraph $(\Theta_{1}\star_{\Psi}\Theta_{2},\gamma)$.
Definition 10.4 (The comb digraph).
Let $A_{1}=\{c,c^{-1}\}$, $A_{2}=\{a,a^{-1}\}$. Let us define the rooted inverse $A_{1}$-digraphs $(\Gamma,z_{0})$ whose underlying digraph is the line
$z_{i}\smash{\stackrel{{\scriptstyle\text{\tiny{$c$}}}}{{\relbar\joinrel\longrightarrow}}}z_{i+1},\;z_{i+1}\smash{\stackrel{{\scriptstyle\text{\tiny{$c^{-1}$}}}}{{\relbar\joinrel\longrightarrow}}}z_{i},\quad\forall i\in\mathbb{Z},$
The other two rooted inverse $A_{2}$-digraphs $(\Lambda_{a},\lambda_{a}),(\Lambda,x_{0})$ are defined as follows. The inverse graph $(\Lambda_{a},\lambda_{a})$ consists of just one vertex $\lambda_{a}$ with loops labelled by $a,a^{-1}$. Instead, $(\Lambda_{a},x_{0})$ is the defined as the line on $A_{2}$:
$x_{i}\smash{\stackrel{{\scriptstyle\text{\tiny{$a$}}}}{{\relbar\joinrel\longrightarrow}}}x_{i+1},\;x_{i+1}\smash{\stackrel{{\scriptstyle\text{\tiny{$a^{-1}$}}}}{{\relbar\joinrel\longrightarrow}}}x_{i},\quad\forall i\in\mathbb{Z},$
@@
By putting $\Theta_{1}=\{(\Gamma,z_{0})\}$ and $\Theta_{2}=\{(\Lambda,x_{0}),(\Lambda_{a},\lambda_{a})\}$, we see that $(\Theta_{1},\Theta_{2},\Psi)$ is a context-free ensemble, thus the free product $(\mathfrak{C},\gamma)=(\Theta_{1}\star_{\Psi}\Theta_{2},z_{0})$ defines a context-free inverse graph called the comb digraph, see Fig. 6. Note that with the convention of the free product that we have given, the root of this inverse graph is $z_{0}$ which is also identified with $x_{0}$. In $\mathfrak{C}$ there are infinite copies of the $c$-line, we denote by $z_{i}^{(j)}$ the $i$-th vertex of the $j$-th copy of this line. Graphically, two vertices $z_{i}^{(j)}$, $z_{i}^{(k)}$ are at the same “$i$-th level”, but they belong to the $j$-th and $k$-th $c$-line, respectively. The $a$-line intersects at the vertices $z_{0}^{(j)}$.
@@
The transition group $G=\mathcal{G}(\mathfrak{C})$ contains a copy of the infinite rank free abelian group $\mathbb{Z}^{\infty}$. In particular, $G\in{\bf CF-TR}$ is not a poly-context-free group.
@@
Consider the set $B=\{g_{n}=c^{n}ac^{-n},n\in\mathbb{Z}\}$. It is not difficult to check that for any vertex $z_{i}^{(j)}$ the generic element $g_{n}^{k}$ acts as follows:
(12) $z_{i}^{(j)}\cdot g_{n}^{k}=\begin{cases}z_{i}^{(j)},\,\,&\mbox{ if }i\neq-n\\ z_{i}^{(j+k)},\,\,&\mbox{ otherwise. }\end{cases}$
for any $j,i,k,n\in\mathbb{Z}$. Now, take $g_{n},g_{m}$ with $m\neq n$. For any vertex $z_{i}^{(j)}$ with $i\neq-n,-m$ we have
$z_{i}^{(j)}\cdot(g_{n}g_{m})=z_{i}^{(j)}=z_{i}^{(j)}\cdot(g_{m}g_{n})$
if $i=-n$, then we have
$z_{i}^{(j)}\cdot(g_{n}g_{m})=z_{i}^{(j+1)}\cdot g_{m}=z_{i}^{(j+1)}=(z_{i}^{(j)}\cdot g_{m})\cdot g_{n}=z_{i}^{(j)}\cdot(g_{m}g_{n})$
and similarly for $i=-m$. Hence, we have $g_{n}g_{m}=g_{m}g_{n}$ in $G$. Moreover, by the action described in (12) we deduce that all the elements in $B$ are distinct and each $g_{n}$ does not have torsion. Therefore, the subgroup $H=\langle B\rangle$ is an abelian subgroup of $G$. To show that $H\simeq\mathbb{Z}^{\infty}$ it is enough to prove that the set $B$ freely generate $H$, that is, there is no relation of the form
$g_{n_{1}}^{k_{1}}\ldots g_{\ell}^{k_{\ell}}=\mathds{1}_{G}$
where the word $g_{n_{1}}^{k_{1}}\ldots g_{\ell}^{k_{\ell}}$ is freely reduced. Let us assume, by contradiction, that such a relation exists. Since $H$ is commutative we may assume without loss of generality that all $k_{i}\neq 0$, and $n_{t}\neq n_{j}$ for all $t,j\in[1,\ell]$ with $t\neq j$. Now, by applying this word to $z_{-n_{1}}^{(j)}$, by (12) we would have $z_{-n_{1}}^{(j)}\cdot(g_{n_{1}}^{k_{1}}\ldots g_{\ell}^{k_{\ell}})=z_{-n_{1}}^{(j+k_{1})}\neq z_{-n_{1}}^{(j)}$ since $k_{1}\neq 0$ and $n_{1}\neq n_{j}$ for $j\neq 1$. However, this is a contradiction. Thus, since each $g_{n}$ is torsion free, we may conclude that $H\simeq\mathbb{Z}^{\infty}$. The last claim of the proposition follows by the fact that a poly-context-free group can not contain $\mathbb{Z}^{\infty}$, as it contains free abelian subgroups of rank $k$ for all $k\in\mathbb{N}$, see [9, Lemma 4.6]. ∎
Figure 7. The context-free inverse graph $\mathfrak{T}$ obtained by taking the free product of the inverse graphs in the upper part, where the gluing maps are defined by taking $\psi^{1}(z_{i})=\lambda_{a}$ for all $z_{i}\neq z_{0}$, and $\psi^{2}(x_{1})=z_{0}$.
We finally address the second point of Question 6.5 regarding the possibility for a transition group to have arbitrarily large finite groups. We prove that such examples may exists by showing a group in CF-TR having an infinite torsion subgroup. The associated graph $\mathfrak{T}$ on the alphabet $A=\{a,c,a^{-1},c^{-1}\}$ is depicted in Fig. 7. Using the notation in Subsection 7 we notice that $\mathfrak{T}$ is locally-quasi-transitive with infinite orbits by taking $n=1$. We have the following proposition.

### v3

@@ 9.1. Context-free graphs and free product
In this section we prove that the free product of directed graphs preserves the property of being context-free. For this purpose, we need to slightly adapt the notion of isomorphism and end-cone of an $A_{1}$-graph to include the notion of vertex coloring. A vertex-coloring of a graph $\Gamma$ is just a map $\chi:V(\Gamma)\to C$ onto some set of colors $C$. An isomorphism $\psi:\Gamma^{(1)}\to\Gamma^{(2)}$ between two $A$-graphs with color maps $\chi_{1}:V(\Gamma)\to C$, $\chi_{2}:V(\Gamma^{(2)})\to C$ is an isomorphism of $A$-graphs that preserves the colorings: $\chi_{1}=\chi_{2}\circ\psi$. For technical reasons that will be clear later, we call a coloring of a rooted $A$-graph $(\Gamma,\gamma_{0})$ any map $\chi:V(\Gamma)\setminus\{\gamma_{0}\}\to C$, and we use the notation $(\Gamma,\gamma_{0};\chi)$ to stress the coloring function. For such colored rooted $A$-graph $(\Gamma,\gamma_{0};\chi)$, we may regard an end-cone $\Gamma(v,\gamma_{0})$ as a colored $A$-graph with coloring map $\chi\mid_{V(\Gamma(v,\gamma_{0}))}$ given by the restriction of $\chi$ to the set of vertices of the end-cone $\Gamma(v,\gamma_{0})$. We therefore say that the colored $A$-graph $(\Gamma,\gamma_{0};\chi)$ is context-free if there are finitely many end-cones up to end-isomorphism (in the aforementioned sense). We stick to the notation introduced in Definition 9.1, and let us suppose that $\Theta_{1}$ ($\Theta_{2}$) is a collection of $A_{1}$-graphs ($A_{2}$-graphs) with $A_{1}\cap A_{2}=\emptyset$ and with gluing maps $\Psi$. We say that $\Gamma^{(j)}\in\Theta_{1}$ ($\Lambda^{(j)}\in\Theta_{2}$) is context-free if the colored rooted $A_{1}$-graph $(\Gamma^{(j)},\gamma_{0}^{j},\psi^{1}_{j})$ ($A_{2}$-graph $(\Lambda^{(j)},\lambda_{0}^{j},\psi^{2}_{j})$) is context-free. If all the colored $A_{i}$-graphs of $\Theta_{1}\cup\Theta_{2}$ are context-free, we say that $(\Theta_{1},\Theta_{2},\Psi)$ is a context-free ensemble. We have the following theorem.
@@
For a context-free ensemble $(\Theta_{1},\Theta_{2},\Psi)$, the free-product $\Theta_{1}\star_{\Psi}\Theta_{2}$ is a context-free $A_{1}\cup A_{2}$-graph.
@@
an then by identifying $z$ with $\psi^{2}_{s}(z)$. We prove that we have finitely many end-cones up to end-isomorphism (in the usual sense). Now, let $\Omega(v^{*}\gamma_{i}^{j}\lambda_{\ell}^{s},\alpha)$ be another end-cone for which there is an end-isomorphism $\varphi:\Lambda^{s}(\lambda^{s}_{t},\lambda^{s}_{0})\to\Lambda^{s}(\lambda^{s}_{\ell},\lambda^{s}_{0})$ of end-cones of the same $A_{2}$-graph $\Lambda^{(s)}$ that preserves the coloring. Since the coloring map $\psi^{2}_{s}$ is commuting with $\varphi$, for any $z\in V(\Lambda^{s}(\lambda^{s}_{t},\lambda^{s}_{0}))$ we have that $\psi^{2}_{s}(z)=\psi^{2}_{s}(\varphi(z))=\gamma_{0}^{h}$, so both $z$ and $\varphi(z)$ have the same color, hence in both the vertex $z$ and $\varphi(z)$ we are gluing the same graph $\Gamma^{(h)}(\uparrow_{\Psi})$. Therefore, it is straightforward to show that $\varphi$ may be extended
@@
by locally extending $\varphi$ from each disjoint copy $\Gamma^{(h)}(\uparrow_{\Psi})$ associated to $z$, to the same graph $\Gamma^{(h)}(\uparrow_{\Psi})$ associated to $\varphi(z)$. Now, since we are performing the identification $z=\psi^{2}_{s}(z)$ in the domain of $\widehat{\varphi}$ and this is the same as identifying $\varphi(z)$ with $\psi^{2}_{s}(\varphi(z))$, we have that $\widehat{\varphi}$ induces an isomorphism $\phi:\Omega(v^{*}\gamma_{i}^{j}\lambda_{t}^{s},\alpha)\to\Omega(v^{*}\gamma_{i}^{j}\lambda_{\ell}^{s},\alpha)$. Note that the frontier vertices of $\Omega(v^{*}\gamma_{i}^{j}\lambda_{t}^{s},\alpha)$ are (up to renaming) the same as $\Lambda^{s}(\lambda^{s}_{t},\lambda^{s}_{0}))$, and similarly for $\Omega(v^{*}\gamma_{i}^{j}\lambda_{\ell}^{s},\alpha)$. Hence, $\phi$ is mapping the frontier vertices of $\Omega(v^{*}\gamma_{i}^{j}\lambda_{t}^{s},\alpha)$ into the frontier vertices of $\Omega(v^{*}\gamma_{i}^{j}\lambda_{\ell}^{s},\alpha)$, hence they are end-isomorphic. We have just shown that the end-isomorphism classes of the end-cones of $\Omega$ of vertices colored by $2$ depends on the end-isomorphism (preserving the coloring) classes of the graphs $\{\Lambda^{(1)},\ldots,\Lambda^{(m)}\}$ of $\Theta_{2}$, which are finitely many. Similarly, one may show the same for the end-cones of the vertices colored by $1$, and this concludes the proof of the theorem. ∎
@@
Figure 5. On the upper part the graphs $\Theta_{1},\Theta_{2}$, in the lower picture the antenna graph $(\Theta_{1}\star_{\Psi}\Theta_{2},\gamma)$.
10. Some examples using the free product of graphs
@@
In this section we use the free product of context-free graphs to construct some examples of interesting CF-TR groups. In the first example, the antenna graph, we show that in general transition groups of context-free graphs are not residually-finite. In the second one, the comb graph, we prove that the associated transition group is $\mathbb{Z}\wr\mathbb{Z}$, so a non-finitely generated group. Hence, it is a solvable co-context-free group that is not poly-context-free, showing that these classes of groups are different and solving the first point of Question 6.5. The third example addresses the second point of Question 6.5: it provides an infinite torsion subgroup, so that the group have arbitrarily large finite subgroups.
Definition 10.1 (The Antenna graph).
Let $A_{1}=\{c,c^{-1}\}$, $A_{2}=\{a,a^{-1},b,b^{-1}\}$. Consider the rooted $A_{1}$-graph $(\Gamma,\gamma)$ which is the line
@@
$\gamma_{i}\smash{\stackrel{{\scriptstyle\raisebox{-2.71246pt}{\tiny$c$}}}{{\relbar\joinrel\longrightarrow}}}\gamma_{i+1},\;\gamma_{i+1}\smash{\stackrel{{\scriptstyle\raisebox{-2.71246pt}{\tiny$c^{-1}$}}}{{\relbar\joinrel\longrightarrow}}}\gamma_{i},\quad\forall i\in\mathbb{Z},\mbox{ where the root }\gamma=\gamma_{0}$
@@
and the two rooted $A_{2}$-graphs $(\Lambda_{a},\lambda_{a}),(\Lambda_{b},\lambda_{b})$ which are lines with loops defined by
@@
$\lambda_{a,i}\smash{\stackrel{{\scriptstyle\raisebox{-2.71246pt}{\tiny$b$}}}{{\relbar\joinrel\longrightarrow}}}\lambda_{a,i},\;\lambda_{a,i}\smash{\stackrel{{\scriptstyle\raisebox{-2.71246pt}{\tiny$b^{-1}$}}}{{\relbar\joinrel\longrightarrow}}}\lambda_{a,i},\,\lambda_{a,i}\smash{\stackrel{{\scriptstyle\raisebox{-2.71246pt}{\tiny$a$}}}{{\relbar\joinrel\longrightarrow}}}\lambda_{a,i+1},\;\lambda_{a,i+1}\smash{\stackrel{{\scriptstyle\raisebox{-2.71246pt}{\tiny$a^{-1}$}}}{{\relbar\joinrel\longrightarrow}}}\lambda_{a,i},\quad\forall i\in\mathbb{Z},\,\,\lambda_{a,0}=\lambda_{a}$
@@
$\lambda_{b,i}\smash{\stackrel{{\scriptstyle\raisebox{-2.71246pt}{\tiny$a$}}}{{\relbar\joinrel\longrightarrow}}}\lambda_{b,i},\;\lambda_{b,i}\smash{\stackrel{{\scriptstyle\raisebox{-2.71246pt}{\tiny$a^{-1}$}}}{{\relbar\joinrel\longrightarrow}}}\lambda_{b,i},\,\lambda_{b,i}\smash{\stackrel{{\scriptstyle\raisebox{-2.71246pt}{\tiny$b$}}}{{\relbar\joinrel\longrightarrow}}}\lambda_{b,i+1},\;\lambda_{b,i+1}\smash{\stackrel{{\scriptstyle\raisebox{-2.71246pt}{\tiny$b^{-1}$}}}{{\relbar\joinrel\longrightarrow}}}\lambda_{b,i},\quad\forall i\in\mathbb{Z},\,\,\lambda_{b,0}=\lambda_{b}$
@@
that are constant maps. By putting $\Theta_{1}=\{(\Gamma,\gamma)\}$ and $\Theta_{2}=\{(\Lambda_{a},\lambda_{a}),(\Lambda_{b},\lambda_{b})\}$, the rooted $A_{1}\cup A_{2}$-graph $(\mathfrak{A},\gamma)=(\Theta_{1}\star_{\Psi}\Theta_{2},\gamma)$ is called the antenna graph, see Fig. 5. Moreover, it is easy to see that $(\Theta_{1},\Theta_{2},\Psi)$ is a context-free-ensemble, thus by Theorem 9.5 $(\mathfrak{A},\gamma)$ is a context-free graph. Note that with the convention of the free product that we have given, the root $\gamma$ of $\Theta_{1}\star_{\Psi}\Theta_{2}$ is identified with $\lambda_{b}$.
@@
[…]
@@
        Now, if $t\neq-k$, then $v^{\prime}c^{-\ell}b^{t}\cdot(b^{k}c^{-\alpha}a^{-s}c^{\alpha}b^{-k}c^{-\alpha})=v^{\prime}c^{-\ell}b^{t+k}c^{-\alpha}\cdot(a^{-s}c^{\alpha}b^{-k}c^{-\alpha})=v^{\prime}c^{-\ell}b^{t}c^{-\alpha}=v$. Otherwise, suppose that t=kt=-k. In this case we have
@@
  • the last case is when $x=a$. In this case we have to subdivide the following two subcases:
@@
The group $G=\mathcal{G}(\mathfrak{A})$ defined by the antenna graph $\mathfrak{A}$ is not residually finite.
@@
In light of Theorem 6.2 and the preceding theorem, we observe that, if Brough Conjecture 1 regarding poly-context-free groups is true, then the class of all poly-context-free groups would be strictly contained within the class CF-TR. Although we are not able to show that the class of poly-context-free is included into our class, in the next example we show that these two classes are different. This is done by defining the following context-free graph.
Figure 6. On the upper part the graphs $\Theta_{1},\Theta_{2}$, in the lower picture the comb graph $(\Theta_{1}\star_{\Psi}\Theta_{2},\gamma)$.
Definition 10.4 (The comb graph).
Let $A_{1}=\{c,c^{-1}\}$, $A_{2}=\{a,a^{-1}\}$. Let us define the rooted $A_{1}$-graphs $(\Gamma,z_{0})$ whose underlying graph is the line
$z_{i}\smash{\stackrel{{\scriptstyle\raisebox{-2.71246pt}{\tiny$c$}}}{{\relbar\joinrel\longrightarrow}}}z_{i+1},\;z_{i+1}\smash{\stackrel{{\scriptstyle\raisebox{-2.71246pt}{\tiny$c^{-1}$}}}{{\relbar\joinrel\longrightarrow}}}z_{i},\quad\forall i\in\mathbb{Z},$
The other two rooted $A_{2}$-graphs $(\Lambda_{a},\lambda_{a}),(\Lambda,x_{0})$ are defined as follows. The graph $(\Lambda_{a},\lambda_{a})$ consists of just one vertex $\lambda_{a}$ with loops labeled by $a,a^{-1}$. Instead, $(\Lambda_{a},x_{0})$ is the defined as the line on $A_{2}$:
$x_{i}\smash{\stackrel{{\scriptstyle\raisebox{-2.71246pt}{\tiny$a$}}}{{\relbar\joinrel\longrightarrow}}}x_{i+1},\;x_{i+1}\smash{\stackrel{{\scriptstyle\raisebox{-2.71246pt}{\tiny$a^{-1}$}}}{{\relbar\joinrel\longrightarrow}}}x_{i},\quad\forall i\in\mathbb{Z},$
@@
By putting $\Theta_{1}=\{(\Gamma,z_{0})\}$ and $\Theta_{2}=\{(\Lambda,x_{0}),(\Lambda_{a},\lambda_{a})\}$, we see that $(\Theta_{1},\Theta_{2},\Psi)$ is a context-free ensemble, thus the free product $(\mathfrak{C},\gamma)=(\Theta_{1}\star_{\Psi}\Theta_{2},z_{0})$ defines a context-free graph called the comb graph, see Fig. 6. Note that with the convention of the free product that we have given, the root of this graph is $z_{0}$ which is also identified with $x_{0}$. In $\mathfrak{C}$ there are infinite copies of the $c$-line, we denote by $z_{i}^{(j)}$ the $i$-th vertex of the $j$-th copy of this line. Graphically, two vertices $z_{i}^{(j)}$, $z_{i}^{(k)}$ are at the same “$i$-th level”, but they belong to the $j$-th and $k$-th $c$-line, respectively. The $a$-line intersects at the vertices $z_{0}^{(j)}$.
@@
The transition group $G=\mathcal{G}(\mathfrak{C})$ is isomorphic to $\mathbb{Z}\wr\mathbb{Z}$. In particular, $G$ is not finitely presented, solvable and contains a copy of the infinite-rank free abelian group $\mathbb{Z}^{\infty}$. Hence it is not poly–context-free.
@@
By Proposition 4.4 we need to find a core-free subgroup of $\mathbb{Z}\wr\mathbb{Z}$ whose corresponding Schreier graph is exactly the graph depicted in Fig. 6. But this figure is exactly Fig. 1 in [21]. ∎
Figure 7. The context-free graph $\mathfrak{T}$ obtained by taking the free product of the graphs in the upper part, where the gluing maps are defined by taking $\psi^{1}(z_{i})=\lambda_{a}$ for all $z_{i}\neq z_{0}$, and $\psi^{2}(x_{1})=z_{0}$.
We finally address the second point of Question 6.5 regarding the possibility for a transition group to have arbitrarily large finite subgroups. We prove that such examples may exists by showing a group in CF-TR having an infinite torsion subgroup. The associated graph $\mathfrak{T}$ on the alphabet $A=\{a,c,a^{-1},c^{-1}\}$ is depicted in Fig. 7. Using the notation in Subsection 8 we notice that $\mathfrak{T}$ is locally-quasi-transitive with infinite orbits by taking $n=1$. We have the following proposition.